# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 3.1 Linear Transformers are RNNs: Attention Through the Lens of Kernels
Reaching the ambitious goal of reducing the cost of autoregressive inference cost (i.e., from $\mathcal{O}(L_{t})$ to $\mathcal{O}(1)$) while attempting to retain performances of vanilla Transformers and still enabling parallel computations (with crucial impact in reduction of the training times in large datasets) is extremely challenging. In Figure 1, the three aforementioned properties are depicted at the bottom, with blue, pink, and yellow boxes, respectively. There exists multiple categories of models that were recently proposed, sharing the intuition of introducing linearity in the computations (notice that this term is used in a two-fold manner: the attention process does not exploit non-linear functions, and the goal is gain linear complexity), sometimes going back to the theory of kernels, to recover stateful models of attention (Figure 5, first column).
@@
Causal Linear Attention.   Interesting properties emerge from the kernel-based formulation when moving to the autoregressive setting. As previously anticipated, this intrinsically requires to implement a form of causal attention, since the future information cannot be accessed. At a generic time $t$, we need to have access to $S_{t}$ and $z_{t}$ (i.e., they are aggregated up to what has been seen so far), and Linear Transformers can be seen as stateful models that, at each time step, update an internal state $X_{t}\vcentcolon=(S_{t},z_{t})$, composed by a recurrent state matrix $S_{t}$ and a recurrent normalization vector $z_{t}$, which are updated iteratively with what is referred to as additive interaction,
@@
with some initial condition $S_{0}$, $z_{0}$ (usually set to zero). Hence, in an autoregressive setting, $y_{t}$ can be computed in a recurrent fashion, multiplying the current query with the $S_{t}$ portion of the state, and normalizing it with its $z_{t}$ portion. This view challenges the traditional distinction between RNNs as “automata-like” constant-size stateful models and Transformers as not being so. Indeed, autoregressive Transformers can be equivalently expressed as an RNN-like sequence processor with $2$-dimensional constant-size states that are updated by sum, as depicted in Figure 6 (right). Noticeably, this recurrent form allows to compress history into the matrix-valued hidden state, thus eliminating the need to maintain and attend over the KV cache during inference. We remark that, when considering the parallel form of causal linear attention, its complexity is still quadratic in $L$. Indeed, due to the presence of causal masking matrix $M$, it is not possible to exploit the associative property of matrix multiplication to reduce the parallel form complexity from quadratic to linear. Noticeably, Linear Transformers combines the best of standard attention and recurrence: at training time (i.e., in non-causal tasks where the full sequence is available in advance), computations can be parallelized and take full advantage of GPUs or other accelerators by leveraging the parallel form. When it comes to inference (e.g., sequential decoding), the cost per time and memory for one prediction is constant, and the internal state $X_{t}=(S_{t},z_{t})$ can be updated at every time step like in RNNs, thanks to their recurrent form. Tasks where vanilla quadratic-time softmax attention cannot be fully-parallelized across time steps can be effectively accelerated, such as in autoregressive decoding, both in the conditional (e.g., machine translation) and unconditional (e.g., sampling from a language model) cases [224]. Additionally, it has been proposed to reduce memory transfer costs by avoiding to materialize states in GPU slow HBM memory (see Section 8.2 for further details).
Random Feature Maps.   Despite the effective computational advantages, Linear Transformers underperform vanilla self-attention, and the main cause has been attributed to the choice of the kernel function [224, 259, 166]. More recent works [40, 41, 224, 259] leverage random feature maps to achieve unbiased estimations of shift-invariant kernels, such as the Gaussian one, in order to approximate effects obtained by using the softmax function in vanilla Transformers. Performer [40, 41] uses random feature maps defined by custom functions $f_{1},\cdots,f_{l}\colon\mathbb{R}\rightarrow\mathbb{R}$ and $h\colon\mathbb{R}^{d_{\text{in}}}\rightarrow\mathbb{R}$. Formally, given an input vector $k_{t}\in\mathbb{R}^{d_{\text{k}}}$ (it could be a key or a query),
@@
where the notation $[a_{i}]_{i=1}^{l}$ indicates a vector obtained by concatenating all the $l$ different $a_{i}$’s, $\omega_{1},\cdots,\omega_{m}$ are random vectors independently drawn from a distribution $\mathcal{D}\in\mathcal{P}(\mathbb{R}^{d_{\text{in}}})$, i.e., $\omega_{1},\cdots,\omega_{m}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{D}$. Choromanski et al. [41] tested trigonometric functions with $h(x)=\exp({\|x\|^{2}/2}),l=2,f_{1}=\sin,f_{2}=\cos$, inspired by random the Fourier feature map [239], that has been proved to be successful in speeding up kernel methods [205] and to approximate softmax [242]. Given that relying on trigonometric functions do not guarantee non-negative attention scores and thus could lead to unstable behaviors, follow-up works [41] proposed positive random feature maps, such as those based on $h(x)=\exp(-\|x\|^{2}/2),l=1,f_{1}=\exp$, guaranteeing unbiased and non-negative approximation of dot-product attention. A concurrent work, Random Feature Attention (RFA) [224], leverages similar intuitions, building a feature map with $h(x)$ set to $1$, with queries and keys $\ell_{2}$-normalized in advance (see [224] for more details), showing benefits with respect to the kernels based on the elu. Additionally, RFA includes a learnable gating mechanism inspired by LSTMs [105] to explicitly model the notion of recency bias and locality, that are not considered in vanilla self-attention. Eq. 10 and 11 are augmented by means of a gating function returning a score $g_{t}$ in $(0,1)$, becoming
@@
where $w_{g}$ is a vector of learnable parameters (there could also a bias term), and $\sigma$ is a sigmoid activation. The multiplicative interaction between the learned scalar gates $10<g_{t}<$ 1 and the hidden state $X_{t}=(S_{t},z_{t})$ exponentially decays past information, favoring more recent contexts. In addition to using random feature maps to approximate standard dot product attention, [224] and [41] explore alternative routes, approximating order-1 arc-cosine kernel. In this case with $h(x)=1,l=1,f_{1}=\mathrm{ReLU}$. This feature map has been show to be effective in various tasks including machine translation and protein sequence modeling.
@@
Fast Weight Programmers and Delta Nets.   Schlag et al. [259] showed the formal equivalence between kernel-based Linear Transformers and the seminal works on Fast Weight Programmers (FWPs) [261, 260]. It turns out that the kernel-based causal perspective of linear attention (Eqs. 10-12 and Figure 6-right) is a FWP with the additional recurrent normalization factor $z_{t}$. The intuition behind the notion of fast weights is to introduce novel dependencies on the weights of the model. A two-network system is proposed, composed of a slow net with slowly-changing weights, which continually reprograms another net with weights that change in a fast manner, making them dependent on the spatio-temporal context of a given input stream. Noticeably, this finding suggests that the recurrent state matrix $S_{t}$ ($\approx W^{(t)}$ in FWP) can be seen as a key-value associative “memory matrix” which gets reprogrammed. ($i$) The “write” operation is obtained by aggregating the results of the outer products of ($\phi$-mapped) keys and values in Eq. 10, also referred to as associations. ($ii$) The “retrieve” operation consists in multiplying the memory matrix by the ($\phi$-mapped) query (see Eq. 12). Schlag et al. [259] argue that endlessly adding new associations to a memory of finite size (Eq. (10)) will inevitably reach a capacity limit. To prevent associations from interfering with each other upon retrieval, keys must be orthogonal, i.e., the feature map size must be large enough to avoid wokring in an overcapacity regime. [^6] With keys embedded in a ${d_{\text{ker}}}$-space, there cannot be more than ${d_{\text{ker}}}$ orthogonal vectors. That is, storing more than ${d_{\text{ker}}}$ associations will result in retrieval issues. Linear attention based on the elu function preserves the dimension of the input key vector (${d_{\text{ker}}}={d_{\text{k}}}$) without modifying the memory capacity, thus, when $L>{d_{\text{ker}}}$, the model might be in such an overcapacity regime. This analysis showcases the limits and sub-optimality of random feature maps, characterized by a $2m$-sized capacity which would require $m$ to go to infinity to yield robust approximations. Interestingly, Schlag et al. introduce a deterministic parameter-free feature map which foster orthogonality in feature space. Specifically, the feature map they propose is such that ${d_{\text{ker}}}=2\nu{d_{\text{k}}}$, and it is defined as
@@
Other Kernel-based Linear Methods.   There exist other works that we can connect to the principles behind linear attention and recurrence. Ecoformer [170] exploits an energy-saving kernel-based hashing (RBF kernel) to map the queries and keys onto low-dimensional binary codes in Hamming space. Kernelized hash functions are learned in a self-supervised manner, driven by the Hamming affinity of the attention scores. CosFormer [233] uses $\phi(x)=\text{ReLu}(x)$ to ensure non-negativity in the attention weights, and a cosine-based re-weighting mechanism to enforce the locality bias in the original softmax attention. The authors of TransNormer [230] identify issues in kernel-based Linear Transformer, both due to unbounded gradients and to attention dilution, i.e., the sparse distribution of attention scores in long sequences. They propose to solve them by normalizing the attention computation and leveraging diagonal attention confined to neighbouring tokens in early layers. In particular, TransNormer leverages a linear kernel, replacing the role of the $z_{t}$ normalization factor of Eq. (12) by Layer Normalization [10] on the attention scores (NormAttention). This results in the most compact form of the update equations of Linear Transformers,
@@ 3.2 Alternative Low-rank Approximations
Kernel-based methods exploited to linearize Transformer self-attention are effective in reducing the computational cost with respect to the sequence length. Still, they result in an complexity quadratic to the model’s feature dimension, $\mathcal{O}(L{d_{\text{k}}}^{2})$, caused by their reliance on matrix dot products, which is unfriendly to large model sizes [313].
@@
where the division is intended to operate element-wise, while $w_{t,i}$ denotes the $(t,i)$-th element of matrix $\mathcal{W}\in\mathbb{R}^{L\times L}$, which is a learnable matrix of pairwise position biases. Indeed, for each input token at position $t$, AFT computes a weighted average of values, the result of which is combined with the query by an element-wise multiplication. This approach eliminates the need for dot product self-attention, even if global interactions between query and values are kept, without requiring to compute and store the attention matrix. As noted by the authors, AFT triggers an implicit attention mechanisms with as many heads as the feature dimensions, wherein the attention matrices have a factorized structure. Such a procedure has a memory complexity which is linear with respect to both the context size and the number of features, making it well-suited for both large inputs and large model sizes, and is generally referred to as channel directed attention. The special configuration in which no position biases are learned, referred to as AFT-simple, can be trivially formalized as,
@@
[…]
where $\mu_{.}$, $W_{.}$, and $p$ are additional trainable parameters. In Eq. (18)-(20), recurrence is implemented by means of a linear interpolation between input at $t$ and $t-1$, referred to as token-shift. In Eq. (21), receptance $r_{t}$ participates in the computation of a gate that acts as a forget gate, which eliminates unnecessary historical information. The recurrent nature of the model goes beyond Eq. (18)-(20), and it also affects Eq. (21), once we consider autoregressive decoding at inference. In fact Eq. (21) can be written in the following recurrent form,
@@
$\hskip 14.22636pt\text{LoRa}(a+(b-a)\odot\mu_{x},A,B,\lambda),$
@@ 3.3 Modeling Recurrence
Amidst the multiple advantages brought by the vanilla self-attention mechanism, early attempts to tackle language modeling with Transformers were slacken by the inability ($i$) to model intra-sequential relations due to the static order dependencies available in standard positional encodings (i.e., the absence of temporal information) and ($ii$) to propagate inter-sequence information among processed contexts. Several attempts to tackle these two issues are presented in the following. Additionally, such approaches inspired recent sub-quadratic methods that split the overall computation into sequence-chunks that are processed in parallel.
Intra-sequence Recurrence Modeling.   An interesting line of work leverages recurrent mechanisms to increase the representational power of features extracted by Transformers, e.g., by injecting locality biases or temporal ordering in the obtained outputs. Chen et al. [37] showed that representations learned by RNN-based encoders can be augmented by the ones learned with a self-attention encoder, resulting in an improvement in performance for RNN-based neural machine translation (NMT) tasks [276]. Inspired by this work, Hao et al. [97] propose to directly model recurrence in an encoder-decoder Transformer architecture for NMT. This is done by leveraging an additional recurrence encoder, ${E_{\text{rec}}(\cdot)}$, that operates in parallel with respect to the standard Transformer encoder, hereinafter referred to as $E(\cdot)$. The authors propose to implement ${E_{\text{rec}}(\cdot)}$ as ($i$) a bidirectional RNN or as ($ii$) an Attentive Recurrent Network (ARN), a model that performs recurrent steps on the features vectors extracted with an attention model from the input representations $U$. For a generic layer $l$, and its input $U^{l}$, ARN computes:
@@
where $F(x_{t-1}^{l},c_{t}^{l})$ denotes an RNN-like transition function (e.g., such as Eq. (2) or the one used in LSTM) with state $x_{t}$, which processes an external input value $c_{t}$. The operation $\text{ATT}(x_{t-1}^{l},U^{l-1})$ is an attention procedure over the layer-input representations $U^{l-1}$, exploiting the previous state $x_{t-1}^{l}$ as query. Outputs from $E(\cdot)$ and $E_{\text{rec}}(\cdot)$ are fused either by a gating mechanism or by sequentially attending them in the decoder. Injecting this kind of recurrent mechanisms into purely attention-based architectures has been proven to boost their performance. Chen et al. [35] attribute this results on the fact that positional embeddings exploited by standard Transformer architectures are based solely on discrete numerical indices (e.g., using trigonometric functions [291]) and are independent of word content. Hence, they are not prone to capture semantic dependencies between words in a sentence. To overcome this issue, the authors in [35] split the embeddings of each input word/token $u_{t}$ into two parts, resulting in two input sequences $U^{p}=(u_{t}^{p})_{t=1}^{L}$ and $U^{r}=(u_{t}^{r})_{t=1}^{L}$. Then, they explicitly learn recurrent positional embeddings (RPEs) on $U_{r}$ with a non-linear RNN, i.e., RPEs are the element of the temporal sequence of states computed by such RNN. At the same time, the other sub-sequence $U^{p}$ is leveraged to compute positional word embeddings following the standard Transformer pipeline. The concatenation of such embeddings and RPE is then given as input to the Transformer encoder (or decoder), equipped with ad-hoc heads to process the recurrent positional information. Leveraging such recurrent embeddings allows to capture order-related dependencies. Huang et al. [111] proposed a block-diagonalization of a linear RNN layer that allows to rewrite it into a lightweight relative positional encoding matrix for a multi-head self-attention, named Recurrent Encoding Matrix (REM). The overall intuition is to encapsulate the recurrent dynamics of the RNN layer into the positional encodings of a multi-head attention layer, leading towards a self-attention with recurrence (RSA). In particular, by considering a linear RNN (obtained from Eq. (2) when $\sigma$ is the identity function), it is possible to write it in the following compact form,
@@
Segment-level Recurrences.   Standard Transformer training is performed on separated fixed-length segments of text, derived from the context window span, without any information flow across such segments, resulting in the inability to capture any longer-term dependency beyond the predefined context. Hence, the ability to temporally connect different contexts becomes extremely important. The basic feature that can overcome such limitations consists in maintaining a cache memory, to be read as an additional input, in order to store the state of previously processed information. This can be easily implemented by exploiting a segment-level recurrence to aggregate the temporal information at a coarser scale, while the token-by-token dependence is learned by self-attention as usual [166]. A seminal work in this direction is Transformer-XL[52], which sequentially process consecutive segments having size $T$, exploiting a segment-level recurrence. Layer-wise representations from previous segments are cached and exploited as an extended context for the current segment. Considering the $l$-th layer of a Transformer architecture and the $s$-th segment, we denote with $U_{s}^{l}$ the $l$-th layer input representation of the $s$-th input segment (i.e., it corresponds to $U_{s}$ when $l=0$ and to the output of the previous Transformer layer for the segment $s$, i.e., $Y_{s}^{l-1}$, when $l>0$). In Transformer-XL, ${U}_{s}^{l}$ is concatenated with the representations from the previous segments, $Y_{s-1}^{l-1}$, to compose a state that is exploited to produce keys and values as follows,

### v2

@@ 3.1 Linear Transformers are RNNs: Attention Through the Lens of Kernels
Reaching the ambitious goal of reducing the cost of autoregressive inference from $\mathcal{O}(L_{t})$ to $\mathcal{O}(1)$ while attempting to retain the performances of vanilla Transformers and still enabling parallel computations (with crucial impact in reducing training times on large datasets) is extremely challenging. In Figure 1, the three aforementioned properties are depicted at the bottom, with blue, pink, and yellow boxes, respectively. There exist multiple categories of models that were recently proposed, sharing the intuition of introducing linearity in the computations (notice that this term is used in a two-fold manner: the attention process does not exploit non-linear functions, and the goal is to gain linear complexity), sometimes going back to the theory of kernels, to recover stateful models of attention (Figure 5, first column).
@@
Causal Linear Attention.   Interesting properties emerge from the kernel-based formulation when moving to the autoregressive setting. As previously anticipated, this intrinsically requires to implement a form of causal attention, since future information cannot be accessed. At a generic time $t$, we need to have access to $S_{t}$ and $z_{t}$ (i.e., they are aggregated up to what has been seen so far), and Linear Transformers can be seen as stateful models that, at each time step, update an internal state $X_{t}\vcentcolon=(S_{t},z_{t})$, composed by a recurrent state matrix $S_{t}$ and a recurrent normalization vector $z_{t}$, which are updated iteratively with what is referred to as additive interaction,
@@
with some initial condition $S_{0}$, $z_{0}$ (usually set to zero). Hence, in an autoregressive setting, $y_{t}$ can be computed in a recurrent fashion, multiplying the current query with the $S_{t}$ portion of the state, and normalizing it with its $z_{t}$ portion. This view challenges the traditional distinction between RNNs as “automata-like” constant-size stateful models and Transformers as not being so. Indeed, autoregressive Transformers can be equivalently expressed as an RNN-like sequence processor with $2$-dimensional constant-size states that are updated by sum, as depicted in Figure 6 (right). Noticeably, this recurrent form allows to compress history into the matrix-valued hidden state, thus eliminating the need to maintain and attend over the KV cache during inference.
We remark that, when considering the parallel form of causal linear attention, its complexity is still quadratic in $L$. Indeed, due to the presence of the causal masking matrix $M$, it is not possible to exploit the associative property of matrix multiplication to reduce the parallel form complexity from quadratic to linear. Noticeably, Linear Transformers combine the best of standard attention and recurrence: at training time (i.e., in non-causal tasks where the full sequence is available in advance), computations can be parallelized and take full advantage of GPUs or other accelerators by leveraging the parallel form. When it comes to inference (e.g., sequential decoding), the cost per time and memory for one prediction is constant, and the internal state $X_{t}=(S_{t},z_{t})$ can be updated at every time step like in RNNs, thanks to their recurrent form. Tasks where vanilla quadratic-time softmax attention cannot be fully parallelized across time steps can be effectively accelerated, such as in autoregressive decoding, both in the conditional (e.g., machine translation) and unconditional (e.g., sampling from a language model) cases [84]. Additionally, it has been proposed to reduce memory transfer costs by avoiding materializing states in GPU slow HBM memory (see Section 8.2 for further details).
Random Feature Maps.   Despite the effective computational advantages, Linear Transformers underperform compared to vanilla self-attention, and the main cause has been attributed to the choice of the kernel function [84, 85, 73]. More recent works [80, 86, 84, 85] leverage random feature maps to achieve unbiased estimations of shift-invariant kernels, such as the Gaussian one, in order to approximate effects obtained by using the softmax function in vanilla Transformers. Performer [80, 86] uses random feature maps defined by custom functions $f_{1},\cdots,f_{l}\colon\mathbb{R}\rightarrow\mathbb{R}$ and $h\colon\mathbb{R}^{d_{\text{in}}}\rightarrow\mathbb{R}$. Formally, given an input vector $k_{t}\in\mathbb{R}^{d_{\text{k}}}$ (it could be a key or a query),
@@
where the notation $a_{i}]_{i=1}^{l}$ indicates a vector obtained by concatenating all the $l$ different $a_{i}$’s, $\omega_{1},\cdots,\omega_{m}$ are random vectors independently drawn from a distribution $\mathcal{D}\in\mathcal{P}(\mathbb{R}^{d_{\text{in}}})$, i.e., $\omega_{1},\cdots,\omega_{m}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{D}$. Choromanski et al. [86] tested trigonometric functions with $h(x)=\exp({\|x\|^{2}/2}),l=2,f_{1}=\sin,f_{2}=\cos$, inspired by the random Fourier feature map [88], that has been proved to be successful in speeding up kernel methods [89] and to approximate softmax [90]. Given that relying on trigonometric functions does not guarantee non-negative attention scores and thus could lead to unstable behaviors, follow-up works [86] proposed positive random feature maps, such as those based on $h(x)=\exp(-\|x\|^{2}/2),l=1,f_{1}=\exp$, guaranteeing unbiased and non-negative approximation of dot-product attention.
A concurrent work, Random Feature Attention (RFA) [84], leverages similar intuitions, building a feature map with $h(x)$ set to $1$, with queries and keys $\ell_{2}$-normalized in advance (see [84] for more details), which shows benefits with respect to the kernels based on the elu. Additionally, RFA includes a learnable gating mechanism inspired by LSTMs [28] to explicitly model the notion of recency bias and locality, which are not considered in vanilla self-attention. Eq. 10 and 11 are augmented by means of a gating function returning a score $g_{t}$ in $(0,1)$, becoming
@@
where $w_{g}$ is a vector of learnable parameters (there could also a bias term), and $\sigma$ is a sigmoid activation. The multiplicative interaction between the learned scalar gates $0<g_{t}<1$ and the hidden state $X_{t}=(S_{t},z_{t})$ exponentially decays past information, favoring more recent contexts. In addition to using random feature maps to approximate standard dot product attention, [84] and [86] explore alternative routes, approximating order-1 arc-cosine kernel. In this case with $h(x)=1,l=1,f_{1}=\mathrm{ReLU}$. This feature map has been show to be effective in various tasks including machine translation and protein sequence modeling.
@@
Fast Weight Programmers and Delta Nets.   Schlag et al. [85] showed the formal equivalence between kernel-based Linear Transformers and the seminal works on Fast Weight Programmers (FWPs) [91, 92]. It turns out that the kernel-based causal perspective of linear attention (Eqs. 10-12 and Figure 6-right) is a FWP with the additional recurrent normalization factor $z_{t}$. The intuition behind the notion of fast weights is to introduce novel dependencies on the weights of the model. A two-network system is proposed, composed of a slow net with slowly-changing weights, which continually reprograms another net with weights that change in a fast manner, making them dependent on the spatio-temporal context of a given input stream. Notably, this finding suggests that the recurrent state matrix $S_{t}$ ($\approx W^{(t)}$ in FWP) can be seen as a key-value associative “memory matrix” which gets reprogrammed. ($i$) The “write” operation is obtained by aggregating the results of the outer products of ($\phi$-mapped) keys and values in Eq. 10, also referred to as associations. ($ii$) The “retrieve” operation consists of multiplying the memory matrix by the ($\phi$-mapped) query (see Eq. 12). Schlag et al. [85] argue that endlessly adding new associations to a memory of finite size (Eq. (10)) will inevitably reach a capacity limit. To prevent associations from interfering with each other upon retrieval, keys must be orthogonal, i.e., the feature map size must be large enough to avoid working in an overcapacity regime. [^7] With keys embedded in a $d_{\text{ker}}$}-space, there cannot be more than $d_{\text{ker}}$} orthogonal vectors. That is, storing more than $d_{\text{ker}}$} associations will result in retrieval issues. Linear attention based on the elu function preserves the dimension of the input key vector (${d_{\text{ker}}}={d_{\text{k}}}$) without modifying the memory capacity, thus, when $L>{d_{\text{ker}}}$, the model might be in such an overcapacity regime. This analysis showcases the limits and sub-optimality of random feature maps, characterized by a $2m$-sized capacity which would require $m$ to go to infinity to yield robust approximations. Interestingly, Schlag et al. introduce a deterministic parameter-free feature map which fosters orthogonality in feature space. Specifically, the feature map they propose is such that ${d_{\text{ker}}}=2\nu{d_{\text{k}}}$, and it is defined as
@@
Other Kernel-based Linear Methods.   There exist other works that we can connect to the principles behind linear attention and recurrence. Ecoformer [101] exploits an energy-saving kernel-based hashing (RBF kernel) to map the queries and keys onto low-dimensional binary codes in Hamming space. Kernelized hash functions are learned in a self-supervised manner, driven by the Hamming affinity of the attention scores. CosFormer [102] uses $\phi(x)=\text{ReLU}(x)$ ensuring non-negativity in the attention weights, and a cosine-based re-weighting mechanism to enforce the locality bias in the original softmax attention. The authors of TransNormer [103] identify issues in kernel-based Linear Transformer, both due to unbounded gradients and to attention dilution, i.e., the sparse distribution of attention scores in long sequences. They propose to solve them by normalizing the attention computation and leveraging diagonal attention confined to neighbouring tokens in early layers. In particular, TransNormer leverages a linear kernel, replacing the role of the $z_{t}$ normalization factor of Eq. (12) by Layer Normalization [104] on the attention scores (NormAttention). This results in the most compact form of the update equations of Linear Transformers,
@@ 3.2 Alternative Low-rank Approximations
Kernel-based methods exploited to linearize Transformer self-attention are effective in reducing the computational cost with respect to the sequence length. However they result in a computational complexity quadratic to the model’s feature dimension, $\mathcal{O}(L{d_{\text{k}}}^{2})$, caused by their reliance on matrix dot products, which is inefficient forlarge model sizes [81].
@@
where the division is intended to operate element-wise, while $w_{t,i}$ denotes the $(t,i)$-th element of matrix $\mathcal{W}\in\mathbb{R}^{L\times L}$, which is a learnable matrix of pairwise position biases. In fact, for each input token at position $t$, AFT computes a weighted average of values, the result of which is combined with the query by an element-wise multiplication. This approach eliminates the need for dot product self-attention, while still preserving global interactions between query and values , and avoids the need to compute and store the attention matrix. As noted by the authors, AFT triggers an implicit attention mechanism with as many heads as the feature dimensions, wherein the attention matrices have a factorized structure. Such a procedure has a memory complexity that is linear with respect to both the context size and the number of features, making it well-suited for both large inputs and large model sizes, and is generally referred to as channel directed attention. The special configuration in which no position biases are learned, referred to as AFT-simple, can be trivially formalized as
@@
$\thinspace r_{t}$ $\thinspace=\thinspace$ $W_{r}(\mu_{r}\odot u_{t}+(1-\mu_{r})\odot u_{t-1}),$ (18)
$\thinspace k_{t}$ $\thinspace=\thinspace$ $W_{k}(\mu_{k}\odot u_{t}+(1-\mu_{k})\odot u_{t-1}),$ (19)
$\thinspace v_{t}$ $\thinspace=\thinspace$ $W_{v}(\mu_{v}\odot u_{t}+(1-\mu_{v})\odot u_{t-1}),$ (20)
$\thinspace y_{t}$ $\thinspace=\thinspace$ $W_{o}\sigma(r_{t})\odot\frac{\sum_{i=1}^{t-1}e^{-(t-i-1)w+k_{i}}\odot v_{i}+e^{p+k_{t}}\odot v_{t}}{\sum_{i=1}^{t-1}e^{-(t-i-1)w+k_{i}}+e^{p+k_{t}}},$ (21)
where $\mu_{.}$, $W_{.}$, and $p$ are additional trainable parameters. In Eq. (18)-(20), recurrence is implemented by means of a linear interpolation between input at $t$ and t1t-1, referred to as token-shift. In Eq. (21), receptance $r_{t}$ participates in the computation of a gate that acts as a forget gate, which eliminates unnecessary historical information. The recurrent nature of the model goes beyond Eq. (18)-(20), and it also affects Eq. (21), once we consider autoregressive decoding at inference. In fact Eq. (21) can be written in the following recurrent form,
@@
$\qquad\text{LoRa}(a+(b-a)\odot\mu_{x},A,B,\lambda),$
@@ 3.3 Modeling Recurrence
Amidst the multiple advantages brought by the vanilla self-attention mechanism, early attempts to tackle language modeling with Transformers were hampered by the inability ($i$) to model intra-sequential relations due to the static order dependencies available in standard positional encodings (i.e., the absence of temporal information), and ($ii$) to propagate inter-sequence information among processed contexts. Several attempts to tackle these two issues are presented in the following. Additionally, such approaches inspired recent sub-quadratic methods that split the overall computation into sequence-chunks that are processed in parallel.
Intra-sequence Recurrence Modeling.   An interesting line of work leverages recurrent mechanisms to increase the representational power of features extracted by Transformers, e.g., by injecting locality biases or temporal ordering in the obtained outputs. Chen et al. [110] showed that representations learned by RNN-based encoders can be augmented by those learned with a self-attention encoder, resulting in an improvement in performance for RNN-based neural machine translation (NMT) tasks [111]. Inspired by this work, Hao et al. [112] propose to directly model recurrence in an encoder-decoder Transformer architecture for NMT. This is done by leveraging an additional recurrence encoder, $E_{\text{rec}}(\cdot)$}, that operates in parallel with respect to the standard Transformer encoder, hereinafter referred to as $E(\cdot)$. The authors propose to implement $E_{\text{rec}}(\cdot)$} as ($i$) a bidirectional RNN or as ($ii$) an Attentive Recurrent Network (ARN), a model that performs recurrent steps on the features vectors extracted with an attention model from the input representations $U$. For a generic layer $l$, and its input $U^{l}$, ARN computes:
@@
where $F(x_{t-1}^{l},c_{t}^{l})$ denotes an RNN-like transition function (e.g., such as Eq. (2) or the one used in LSTM) with state $x_{t}$, which processes an external input value $c_{t}$. The operation $\text{ATT}(x_{t-1}^{l},U^{l-1})$ is an attention procedure over the layer-input representations $U^{l-1}$, exploiting the previous state $x_{t-1}^{l}$ as query. Outputs from $E(\cdot)$ and $E_{\text{rec}}(\cdot)$ are fused either by a gating mechanism or by sequentially attending them in the decoder. Injecting this kind of recurrent mechanisms into purely attention-based architectures has been proven to boost their performance. Chen et al. [113] attribute these results on the fact that positional embeddings exploited by standard Transformer architectures are based solely on discrete numerical indices (e.g., using trigonometric functions [26]) and are independent of word content. Hence, they are not prone to capture semantic dependencies between words in a sentence. To overcome this issue, the authors in [113] split the embeddings of each input word/token $u_{t}$ into two parts, resulting in two input sequences $U^{p}=(u_{t}^{p})_{t=1}^{L}$ and $U^{r}=(u_{t}^{r})_{t=1}^{L}$. Then, they explicitly learn recurrent positional embeddings (RPEs) on $U_{r}$ with a non-linear RNN, i.e., RPEs are the elements of the temporal sequence of states computed by such RNN. At the same time, the other sub-sequence $U^{p}$ is leveraged to compute positional word embeddings following the standard Transformer pipeline. The concatenation of such embeddings and RPE is then given as input to the Transformer encoder (or decoder), equipped with ad-hoc heads to process the recurrent positional information. Leveraging such recurrent embeddings allows the model to capture order-related dependencies.
Huang et al. [114] proposed a block-diagonalization of a linear RNN layer that allows to rewrite it into a lightweight relative positional encoding matrix for a multi-head self-attention, named Recurrent Encoding Matrix (REM). The overall intuition is to encapsulate the recurrent dynamics of the RNN layer into the positional encodings of a multi-head attention layer, leading towards a self-attention with recurrence (RSA). In particular, by considering a linear RNN (obtained from Eq. (2) when $\sigma$ is the identity function), it is possible to write it in the following compact form,
@@
Segment-level Recurrences.   Standard Transformer training is performed on separate fixed-length segments of text, derived from the context window span, without any information flow across such segments, resulting in the inability to capture any longer-term dependency beyond the predefined context. Hence, the ability to temporally connect different contexts becomes extremely important. The basic feature that can overcome such limitations consists of maintaining a cache memory, to be read as an additional input, in order to store the state of previously processed information. This can be easily implemented by exploiting a segment-level recurrence to aggregate the temporal information at a coarser scale, while the token-by-token dependencies are learned by self-attention as usual [73]. A seminal work in this direction is Transformer-XL[117], which sequentially processes consecutive segments having size $T$, exploiting a segment-level recurrence. Layer-wise representations from previous segments are cached and exploited as an extended context for the current segment. Considering the $l$-th layer of a Transformer architecture and the $s$-th segment, we denote with $U_{s}^{l}$ the $l$-th layer input representation of the $s$-th input segment (i.e., it corresponds to $U_{s}$ when $l=0$ and to the output of the previous Transformer layer for the segment $s$, i.e., $Y_{s}^{l-1}$, when $l>0$). In Transformer-XL, ${U}_{s}^{l}$ is concatenated with the representations from the previous segments, $Y_{s-1}^{l-1}$, to compose a state that is exploited to produce keys and values as follows,